(3x+1)dx / ((x+1)^12(3x-1)^12)


1
Expert's answer
2021-02-24T12:40:27-0500

"\large\frac{(3x+1)dx}{((x+1)^{12}(3x-1)^{12})}" Let's solve the problem f(x)dx is meaning of integral


"\int\large\frac{(3x+1)dx}{((x+1)^{12}(3x-1)^{12})}"

Substitute "u = (x+1)(3x-1) \to \large\frac{du}{dx}""=3(x+1)+(3x-1) \to dx = \large\frac{du}{3(x+1)+3x-1}"

"\to dx= \large\frac{du}{2(3x+1)}"

"\large\frac{1}{2}\int\frac{du}{u^{12}}" "\to(Apply \space power \space rule \space \int u^ndu = \large\frac{u^{n+1}}{n+1})"  with n = -12

                "= \large\frac{1}{2} * (\frac{-1}{11u^{11}})" "=\large\frac{-1}{22u^{11}}"

Undo substitution "u = (x+1)(3x-1)" "\to \large\frac{-1}{22(x+1)^{11}(3x-1)^{11}}"

"\large\int\large\frac{3x+1}{(x+1)^{12}(3x-1)^{12}}dx = -\large\frac{1}{22(x+1)^{11}(3x-1)^{11}}+C"